OpenStudy (anonymous):

horizontal line that passes through the point (-5, 9).

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

what u need to find of the line?

OpenStudy (anonymous):

i need to find a horizontal line that passes through (-5,9)

OpenStudy (uri):

Do you have options? I think it's -5

OpenStudy (anonymous):

i think its negative 5 but idnt know if its y or x

OpenStudy (anonymous):

Slope of horizontal line m=0 thus eq of line passing through (-5, 9) in slope point form is given as y-y1=m(x-x1) y-9=0(x+5) i.e. y-9 =0 i.e. y=9 Hence the required eq of the given line is y=9

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

So the line passes through the point (0,9)

OpenStudy (uri):

I think it's y=-5 anyway :p